OpenStudy (anonymous):

find the complex zeros of f(x)=x^3-5x^2+12x-8. Write f in factored form.

OpenStudy (anonymous):

First you factor it as (-1 + x) (8 - 4 x + x^2), the roots are x=1 and x = 2 - 2 I and x= 2+ 2 i You can practice factoring polynomials on http://www.saab.org/mathdrills/factor.cgi

OpenStudy (campbell_st):

well f(1) = 0 so (x -1) is a factor do the polynomial division to get the quadratic factor then solve the quadratic

OpenStudy (anonymous):

thanks @eliassaab